OpenStudy (lisa2345):

Why does my question look like this? |-4+8| & -|-3+5|

myininaya (myininaya):

why does it look like that? do you mean it to look some other way?

OpenStudy (lisa2345):

No that's how the question was asked with these "|" on each side. I'm wondering what kind of question it is.

OpenStudy (lisa2345):

@myininaya

myininaya (myininaya):

so not why you mean you need help on the question

myininaya (myininaya):

ok well first do you know how to add -4 and 8?

OpenStudy (lisa2345):

it's 12

myininaya (myininaya):

-4+8=4 8-4=4 -4-8=-12 -8-4=-12 4+8=12 8+4=12 4-8=-4 -8+4=-4

myininaya (myininaya):

-4+8 is 4 because we have to subtract 8 and 4 because one of them has a - while the one has a + 8-4=4

myininaya (myininaya):

|4| means you need to calculate the distance that 4 is from 0 how many units is 4 from 0?

OpenStudy (lisa2345):

3?

myininaya (myininaya):